# Conceptualization; Methodology; Resources; Investigation; Writing original draft; Writing, review, and editing

### v1

@@ 4.6 Other
5 Open issues and future directions
@@
Transformers-based CI techniques: Transformer-based techniques, such as connectionist temporal classification (CTC), bidirectional encoder representations from transformer (BERT), and others, proved in the literature that have the potential to greatly enhance the functioning of ASR [107, 108]. By leveraging the self-attention mechanism, transformers can improve speech intelligibility by effectively suppressing background noise and modeling long-range dependencies. They can also aid in acoustic scene analysis, separating and prioritizing important auditory information in complex environments. Transformers can build language models that enhance ASR systems, improving speech comprehension for users. Additionally, transformers enable personalized sound processing by adapting stimulation patterns and processing parameters based on user-specific preferences. They facilitate multi-modal integration, combining audio and visual inputs to enhance speech recognition and sound localization. Furthermore, transformers support long-term learning and adaptation, continually optimizing CI performance over time. These advancements offer promising prospects for improving auditory experiences and overall quality of life for CI users.
Exploring chat-bots-based CI capabilities: Chat-bot techniques offer several opportunities to enhance the functioning of CIs. They can provide real-time support, troubleshooting, and personalized rehabilitation programs for users, empowering them to address common issues and improve their auditory skills. Chat-bots enable remote monitoring, allowing users to share data and receive adjustments to their device settings without in-person appointments. They also offer emotional and psychological support, fostering a sense of community and well-being. Chat-bots contribute to data collection for research and development, aiding in the improvement of CI technology and rehabilitation protocols. Additionally, chat-bots employML to continuously learn from user interactions, improving their responses and understanding over time. These techniques have the potential to enhance the overall user experience, outcomes, and accessibility of CI services.
6 Conclusion
@@
Acknowledgement
The second author acknowledges that the study was partially funded by the Algerian Ministry of Higher Education and Scientific Research (Grant No. PRFU–A25N01UN260120230001).

### v3

@@ IV-F Other
V Open issues and future directions
@@
Transformers-based CI techniques: AI researchers have adopted the CNN-LSTM model that excels in capturing spatial and temporal features, enhancing performance in sequential data tasks [8, 122, 123]. However, Transformer-based ASR techniques, such as connectionist temporal classification (CTC), bidirectional encoder representations from transformer (BERT), and others, proved in the literature that have the potential to greatly enhance the functioning of ASR [124, 125, 126]. By leveraging the self-attention mechanism, transformers can improve speech intelligibility by effectively suppressing background noise and modeling long-range dependencies. They can also aid in acoustic scene analysis, separating and prioritizing important auditory information in complex environments. Transformers can build language models that enhance ASR systems, improving speech comprehension for users. Additionally, transformers enable personalized sound processing by adapting stimulation patterns and processing parameters based on user-specific preferences. They facilitate multi-modal integration, combining audio and visual inputs to enhance speech recognition and sound localization. Furthermore, transformers support long-term learning and adaptation, continually optimizing CI performance over time. These advancements offer promising prospects for improving auditory experiences and overall quality of life for CI users.
Exploring chat-bots-based CI capabilities: Chat-bot techniques offer several opportunities to enhance the functioning of CIs. They can provide real-time support, troubleshooting, and personalized rehabilitation programs for users, empowering them to address common issues and improve their auditory skills. Chat-bots enable remote monitoring, allowing users to share data and receive adjustments to their device settings without in-person appointments. They also offer emotional and psychological support, fostering a sense of community and well-being. Chat-bots contribute to data collection for research and development, aiding in the improvement of CI technology and rehabilitation protocols. Additionally, chat-bots employML to continuously learn from user interactions, improving their responses and understanding over time. These techniques have the potential to enhance the overall user experience, outcomes, and accessibility of CI services. For example, in [127], the effectiveness of ChatGPT-4 in providing postoperative care information to CI patients was evaluated. Five common questions were posed to ChatGPT-4, and its responses were analyzed for accuracy, response time, clarity, and relevance. The results showed that ChatGPT-4 provided accurate and timely responses, making it a reliable supplementary resource for patients in need of information.
VI Conclusion
@@
Data availability
Data will be made available on request.
Conflict of Interest
The authors declare no conflicts of interest.